OpenStudy (anonymous):

when 3 times a number is subtracted from 39, the result is 3 plus the number. what is the number?

OpenStudy (anonymous):

3x-39=3+x ==> x=21

OpenStudy (anonymous):

I mean 21/2

OpenStudy (anonymous):

39-3x=3+x

OpenStudy (anonymous):

4x=36 x=9

OpenStudy (anonymous):

this is the equation: 3x-39=3+x 3x-x=3+39 2x=42 *divided by 2 X=21

OpenStudy (anonymous):

39 - 3x = 3 + x 39-3=x+3x 36=4x 9=x